# SCott: Accelerating Diffusion Models with Stochastic Consistency Distillation

### v1

@@ 3.2 SDE Solvers for Diffusion Models
$\small f_{t}:=\frac{\mathrm{dlog}\alpha_{t}}{\mathrm{d}t},\quad g^{2}_{t}:=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{dlog}\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}.$ (4)
We can discretize Equation 3 over time to get an approximated solution with various SDE solvers (Gonzalez et al., 2023; Cui et al., 2023; Zhang & Chen, 2022).
Benefits of SDE solvers. According to the theoretical analysis of (Xu et al., 2023a), the divergence between $p({\bm{z}}_{0})$ and the sample distribution $p(\hat{{\bm{z}}}_{0})$ stems from the discretization errors along the sampling trajectory and the approximation error between the model and the ground-truth score function (Song et al., 2020c). When the number of function evaluations (NFE) is small, SDE solvers exhibit larger discretization errors than ODE ones—$O(\delta^{2})$ v.s. $O(\delta^{\frac{3}{2}})$ with $\delta$ as the step size for discretization. On the contrary, the discretization errors become less significant as $\delta$ shrinks and the approximation errors dominate, thus SDE solvers achieve higher sample quality than ODE ones thanks to the injection of noise for correcting approximation errors (Karras et al., 2022).
@@
We validate these arguments with results in Table 1, where the SDE solver ER-SDE 5 (Cui et al., 2023) outperforms ODE solvers DDIM and DPM++ (Lu et al., 2022b) in terms of both sample quality and diversity with 50 sampling steps. With such observations, we are motivated to explore the possibility of combining SDE solvers and CD, aiming at acquiring stronger teachers for CD.
4 Methodology
@@
Our goal is to generate high-quality samples within a few-step inference while maintaining the trade-off to improve sample quality with additional sampling steps. Previous methods are either failed to generate high-quality samples within $1-4$ steps (Lu et al., 2022a; Luo et al., 2023a; Salimans & Ho, 2022) or unable to trade additional sampling steps for improved outcomes (Liu et al., 2023; Xu et al., 2023b; Sauer et al., 2023b). We observe that CMs enjoy the trade-off by alternating denoising and noise injection at inference time. However, CD approaches haven’t unleashed the potential power of the pre-trained teacher DMs, considering ODE solvers underperform SDE ones when the NFE is large. We first lay out a justification for the feasibility of combining CD with SDE solvers. Then, we identify several critical factors to make SCott workable for high-resolution text-to-image generation. We also introduce GAN loss to further enhance the sampling quality at rare steps. Our method is summarized in Figure 2.
@@
Figure 2: Overview of SCott. SCott distills a pre-trained teacher DM into a student one for accelerated sampling. Compared to the vanilla consistency distillation approach, we introduce a multi-step SDE solver to establish a stronger and more versatile teacher. We refine the student model to yield a Gaussian distribution to account for the uncertainty in SDE solving. We also include an adversarial learning loss to boost the sample quality with rare sampling steps. Note that we omit the EMA module for brevity.
4.1 Justification of Using SDE Solvers for CD
@@
Despite our motivation for combining SDE solvers and CD being sensible from the perspective of model distillation, a major concern arises as to whether it is reasonable to do so because CD is originally defined on ODE trajectories. This section provides some empirical investigation on this.
Specifically, we experiment on a simple 1D task—characterizing a Gaussian mixture distribution with 3 uniform components $p_{\rm{data}}({\bm{x}}):=\frac{1}{3}\mathcal{N}({\bm{x}};-1.5,0.2^{2}\mathbf{I})+\frac{1}{3}\mathcal{N}({\bm{x}};0,0.2^{2}\mathbf{I})+\frac{1}{3}\mathcal{N}({\bm{x}};1.5,0.2^{2}\mathbf{I})$. We first train a DM, instantiated as a 4-layer MLP, with a standard Gaussian prior on it, and then perform CD based on Equation 1 with SDE and ODE solvers respectively. In particular, we select the EDM solver (Karras et al., 2022) with the noise coefficient equaling 0 as the ODE solver and the EDM solver with that equaling 2 as the SDE solver.
We plot the one-step sampling results of the trained models in Figure 3, where the Gaussian noise and sampling trajectory are also displayed. As shown, the stochastic CD can successfully map noise points to suitable target samples, even with higher accuracy than the vanilla CD. One possible explanation for this phenomenon is that, for a given $t_{m}$, the sampled states from the SDE solver are likely to surround the deterministic states sampled from the ODE solver. This situation can be viewed as a form of data augmentation, which helps the student model better align and rectify its predictions. These results provide concrete evidence that we can indeed use SDE solvers for CD.
4.2 Stochastic Consistency Distillation (SCott)
@@
Controlling the level of noise in SDE solvers. It is a natural idea to control the level of random noise injected in SDE solvers to stabilize training. Taking the DDIM solver in Equation 2 for example, the $\sigma_{t_{n}}$ controls the intensity of injected noise and typically we can introduce a coefficient $\eta$ for scaling it (Song et al., 2020a):
$\small\sigma_{t_{n}}(\eta):=\eta\sqrt{(1-\alpha_{t_{n-1}})/(1-\alpha_{t_{n}})}\sqrt{1-\alpha_{t_{n}}/\alpha_{t_{n-1}}}.$ (5)
Increasing $\eta$ from 0 to a positive value results in a set of SDE solvers with increasingly intensive noise, starting from an ODE one. We empirically experiment with the solvers for SCott and present the results in Table 4. As shown, we indeed need to select a reasonable noise level to enjoy the benefits of SDE solvers for error correction while avoiding introducing excessive variance. Following this insight, we include the more advanced ER-SDE solver (Cui et al., 2023) to SCott. The novel noise scaling function in the ER-SDE solver also provides a simple way to control the intensity of noise, which aligns with our requirements for SDE solvers.
Multi-step sampling. The discretization errors of SDE solvers are larger than those of ODE solvers (Xu et al., 2023a). A simple remediation to this is to decrease the step size, but doing so leads to slow convergence and degraded results for SCott, consistent with the results in (Luo et al., 2023a). To address this, we propose a multi-step sampling strategy for SCott. Specifically, for a sample ${\bm{z}}_{t_{n}}$ at the time $t_{n}$ on the SDE trajectory, to obtain the estimated state $\hat{{\bm{z}}}_{t_{m}}$ for a preceding timestep $t_{m}$, we split the time interval $t_{n}-t_{m}$ into $h$ intervals, namely, sampling with a step size of $\frac{t_{n}-t_{m}}{h}$. We denote the solution as $\hat{{\bm{z}}}_{t_{m}}=\Phi_{SDE}({\bm{\epsilon}}_{\theta},{\bm{z}}_{t_{n}},t_{n},t_{m},h)$.
@@
A Gaussian student and the parameterization. To accommodate the uncertainty in stochastic consistency distillation, we introduce a Gaussian formula to the student model and CD objective. For ${\bm{z}}_{t_{n}}$, we predict both the Gaussian mean and variance of the sample. Specifically, we reparameterize our network ${\bm{f}}_{{\bm{\theta}}}$ to produce $[{\bm{\mu}}_{t_{n}},{\bm{\sigma}}^{2}_{t_{n}}]={\bm{f}}_{{\bm{\theta}}}({\bm{z}}_{t_{n}},{\bm{c}},\omega,t_{n})$ where we also incorporate the condition ${\bm{c}}$ and the classifier-free guidance (CFG) (Ho & Salimans, 2022) scale $\omega$ as model input. In practice, we only double the output channels of the final layer of the model to fulfill the need to produce both the mean and variance. Then, we extend the original $\ell_{2}$ loss for CD as the Kullback–Leibler (KL) divergence between Gaussians:
$\small\begin{split}&\mathcal{L}_{\mathcal{KL}}({\bm{\theta}},{\bm{\theta}}^{-};\Phi_{SDE}):=\\ &\quad\quad\mathbb{E}_{t_{n},{\bm{z}}_{t_{n}}}\Big{[}\lambda(t_{n})D_{\mathrm{KL}}(\mathcal{N}({\bm{\mu}}_{t_{n}},{\bm{\sigma}}^{2}_{t_{n}})\|\mathcal{N}(\hat{{\bm{\mu}}}_{t_{m}},\hat{{\bm{\sigma}}}^{2}_{t_{m}}))\Big{]}\end{split}$ (6)
where $[\hat{{\bm{\mu}}}_{t_{m}},\hat{{\bm{\sigma}}}^{2}_{t_{m}}]={\bm{f}}_{{\bm{\theta}}-}(\hat{{\bm{z}}}_{t_{m}},{\bm{c}},\omega,t_{m})$. Gaussian formulation allows for a more reliable representation of the consistency target, capturing not only its expectation but also the uncertainty. We observe that such uncertainty can lead to an increase in diversity among generated samples.
Algorithm 1 Stochastic Consistency Distillation
  Input: dataset $\mathcal{D}$, initial consistency model parameter ${\bm{\theta}}$, discriminator parameter ${\bm{\phi}}$, SDE solver $\Phi_{SDE}$, encoder $E(\cdot)$, noise schedule $\alpha_{t},\sigma_{t}$, sampling steps $h$, adversarial loss $\mathcal{L}_{adv}(\cdot,\cdot)$, weight $\lambda(\cdot)$, CFG scale $\omega$, EMA rate $\mu$, loss coefficient $\lambda_{adv}$ ${\bm{\theta}}^{-}\leftarrow{\bm{\theta}}$
  repeat
     Sample $({\bm{x}}_{0},{\bm{c}})\sim\mathcal{D}$ Sample ${\bm{z}}_{t_{n}}\sim\mathcal{N}(\alpha_{t_{n}}E({\bm{x}}_{0}),\sigma^{2}_{t_{n}}\mathbf{I})$ $\hat{{\bm{z}}}_{t_{m}}\leftarrow\Phi_{SDE}({\bm{\epsilon}}_{\theta},{\bm{z}}_{t_{n}},t_{n},t_{m},h)$ Calculate $\mathcal{L_{KL}}({\bm{\theta}},{\bm{\theta}}^{-};\Phi_{SDE})$ with Equation 6 Calculate $\mathcal{L}_{adv}({\bm{\theta}},{\bm{\phi}})$ with Equation 7 Calculate $\mathcal{L}_{SCott}({\bm{\theta}},{\bm{\phi}})$ with Equation 8 ${\bm{\theta}}\leftarrow{\bm{\theta}}-\frac{\partial}{\partial{\bm{\theta}}}\mathcal{L}_{SCott}({\bm{\theta}},{\bm{\phi}})$ ${\bm{\phi}}\leftarrow{\bm{\phi}}-\frac{\partial}{\partial{\bm{\phi}}}\mathcal{L}_{SCott}({\bm{\theta}},{\bm{\phi}})$ ${\bm{\theta}}^{-}\leftarrow$ stop_grad$(\mu{\bm{\theta}}^{-}+(1-\mu){\bm{\theta}})$
  until convergence
4.3 Introducing GAN Loss into SCott
Our SCott distills the pre-trained teacher to obtain the accelerated DM ${\bm{f}}_{\bm{\theta}}$, so its performance is upper bounded by the teacher. Suggested by CTM (Kim et al., 2023) which leverages GAN to improve trajectory estimation in the process of distillation, we integrate GAN into SCott to mitigate the teacher’s restriction. This renders SCott to produce more realistic outputs, particularly at rare steps. Since SCott predicts the clean image $\hat{{\bm{z}}}_{0}$, we can directly perform adversarial learning between real image ${\bm{z}}_{0}$ and generated image $\hat{{\bm{z}}}_{0}$. However, we find such a naive design makes training unstable for text-to-image tasks. To surmount this challenge, the timestep $t_{n}$ and text ${\bm{c}}$ are propagated into the discriminator as conditions.
@@ 4.3 Introducing GAN Loss into SCott
To accelerate training, the discriminator is initialized by pre-trained U-Net in SD (Rombach et al., 2022). However, training the whole U-Net is expensive. Therefore, we freeze the encoder of the U-Net and only train the low-rank adaptation (LoRA) (Hu et al., 2021) parameters of the decoder of the network. By employing this approach, only $3.64\%$ of the parameters in the U-Net are updated, which not only contributes to computational efficiency but also increases the overall stability of the adversarial training.
Following (Sauer et al., 2023a), we train our discriminator $D_{{\bm{\phi}}}$ with hinge loss:
$\small\begin{split}&\mathcal{L}_{adv}({\bm{\theta}},{\bm{\phi}})=\mathbb{E}_{{\bm{z}}_{0},{\bm{c}}}\Big{[}\max(0,1-D_{{\bm{\phi}}}({\bm{z}}_{0},{\bm{c}},0))\Big{]}\\ &\quad+\mathbb{E}_{{\bm{z}}_{t_{n}},{\bm{c}},t_{n}}\Big{[}\max(0,1+D_{{\bm{\phi}}}(\bar{{\bm{f}}}_{{\bm{\theta}}}({\bm{z}}_{t_{n}},{\bm{c}},\omega,t_{n}),{\bm{c}},t_{n}))\Big{]},\end{split}$ (7)
where $\bar{{\bm{f}}}_{{\bm{\theta}}}$ denotes the function that produces the Gaussian mean within the model ${{\bm{f}}}_{{\bm{\theta}}}$.
@@
Overall, SCott is trained with the following objective:
$\small\begin{split}\mathcal{L}_{SCott}({\bm{\theta}},{\bm{\phi}})=\mathcal{L}_{\mathcal{KL}}({\bm{\theta}},{\bm{\theta}}^{-};\Phi_{SDE})+\lambda_{adv}\mathcal{L}_{adv}({\bm{\theta}},{\bm{\phi}}).\end{split}$ (8)
@@
In practice, we set $\lambda_{adv}=0.4$ to control the strength of the discriminator for refining the outputs of ${\bm{f}}_{\bm{\theta}}$. To save GPU memory, the training is performed in latent space. We present the pseudo-code for SCott in 4.
5 Experiments
In this section, we elaborate on the experimental results of our proposed SCott model for the text-to-image generation task. We start with the comparison with other works in Section 5.1, and then ablate components of SCott in Section 5.2, highlighting the effectiveness of our proposed components.
5.1 Comparison with Other Methods
@@
Table 2: Comparisons with the state-of-the-art methods on MSCOCO-2017 5K in terms of FID, CS, CR. All models are based on SD1.5.
Method Step Time (s) FID $\downarrow$ CS $\uparrow$ CR $\uparrow$
DPM++ (Lu et al., 2022b) 25 0.88 20.1 0.318 0.9118
DDIM (Song et al., 2020b) 50 -- 20.3 0.318 0.9130
PD (Salimans & Ho, 2022) 4 0.21 26.4 0.300 --
CAD (Meng et al., 2023) 8 0.34 24.2 0.300 --
LCM (Luo et al., 2023a) 2 0.13 30.4 0.293 0.7882
INSTAFLOW-0.9B (Liu et al., 2023) 1 0.09 23.4 0.304 --
INSTAFLOW-1.7B (Liu et al., 2023) 1 0.12 22.4 0.309 --
UFOGEN (Xu et al., 2023b) 1 0.09 22.5 0.311 --
UFOGEN (Xu et al., 2023b) 4 -- 22.1 0.307 --
SCott (Ours) 1 0.09 26.5 0.295 0.8442
SCott (Ours) 2 0.13 22.1 0.309 0.9169
SCott (Ours) 4 0.21 21.9 0.312 0.9210
Comparison on MSCOCO-2017 5K with SD1.5. To kick-start the comparisons with the state-of-the-art methods, we evaluate the MSCOCO-2017 5K validation dataset (Lin et al., 2014). Zero-shot FID and CLIP Score (CS) with ViT-g/14 backbone (Radford et al., 2021) are exploited as objective metrics. To measure diversity, Coverage (CR) (Naeem et al., 2020) is used as another metric. Table 2 summarizes the performance of our SCott and comparative methods consisting of PD (Salimans & Ho, 2022), CAD (Meng et al., 2023), LCM (Luo et al., 2023a), InstaFlow (Liu et al., 2023), UFOGen (Xu et al., 2023b). Since PD (Salimans & Ho, 2022), CAD (Meng et al., 2023), InstaFlow (Liu et al., 2023), and UFOGen (Xu et al., 2023b) do not list CR value, we leave them --. Time denotes inference time on a single A100. LCM is implemented according to the official in our setting.
Among all the methods, our 2-step Scott presents superior FID and CS values than 4-step PD (Salimans & Ho, 2022), 8-step CAD (Meng et al., 2023) 1-step InstaFlow-0.9B (Liu et al., 2023), 2-step LCM (Luo et al., 2023a). It is impressive to see that our 2-step SCott beats InstaFLow-1.7B which doubles the parameter size. These results demonstrate our SCott significantly enhances the quality of the generated images while reducing inference steps. The improvements lie in that the proposed $\mathcal{L_{KL}}$ in SCott is able to remarkably reduce the sampling steps and adversarial loss further improves image quality. Interestingly, the CS decreases by 0.004 when the inference step increases from 1 to 4 for UFOGen (Xu et al., 2023b). Meanwhile, by changing the inference step from 1 to 4, all the metrics are improved for our SCott, and 4-step Scott also outperforms 4-step UFOGen regarding both FID and CS. The result indicates the improvement is limited for UFOGen when increasing sampling steps while our method is much more powerful in enhancing image quality with increasing steps, which is meaningful in the scenarios requiring high-quality images with more affordable computational budgets. The proposed SCott obtains higher CR than LCM (Luo et al., 2023a), and even outperforms the ODE solvers, 25-step DPM++ (Lu et al., 2022b) and 50-step DDIM (Song et al., 2020b), indicating our SCott successfully increases diversity due to the introduced randomness.
Figure 5 presents the qualitative comparisons involving InstaFlow and LCM. Notably, 2-step SCott gains significant improvements over 2-step LCM and InstaFlow in terms of image quality and text-to-image alignment. We can also observe that our generated images exhibit sharper textures and finer details, compared to the images generated by InstaFlow and LCM.

### v4

@@ Introduction
$f_{t}:=\frac{\mathrm{dlog}\alpha_{t}}{\mathrm{d}t},\quad g^{2}_{t}:=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{dlog}\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}.$ (4)
We can discretize Equation 3 over time to get an approximated solution with various SDE solvers (Gonzalez et al. 2023; Cui et al. 2023; Zhang and Chen 2022).
Benefits of SDE Solvers. According to the theoretical analysis of (Xu et al. 2023), the divergence between $p({\bm{z}}_{0})$ and the sample distribution $p(\hat{{\bm{z}}}_{0})$ stems from the discretization errors along the sampling trajectory and the approximation error between the model and the ground-truth score function (Song et al. 2020). When the number of function evaluations (NFE) is small, SDE solvers exhibit larger discretization errors than ODE ones—$O(\delta^{\frac{3}{2}})$ v.s. $O(\delta^{2})$ with $\delta$ as the step size for discretization. On the contrary, the discretization errors become less significant as $\delta$ shrinks and the approximation errors dominate, thus SDE solvers achieve higher sample quality than ODE ones thanks to the injection of noise for correcting previous approximation errors (Karras et al. 2022).
@@
We validate these arguments with results in Table 1, where the ER-SDE 5 solver (Cui et al. 2023) outperforms ODE solvers DDIM and DPM++ (Lu et al. 2022b) in terms of both sample quality and diversity with 50 sampling steps. Besides, the noise injected in SDE solvers implicitly makes stochastic data augmentation for CD. With such observations, we are motivated to explore the possibility of combining SDE solvers and CD, aiming at acquiring stronger teachers for CD.
Methodology
@@
Previous diffusion acceleration methods either fail to generate high-quality samples within $4$ steps (Lu et al. 2022a; Luo et al. 2023a) or are limited to one-step generation, unable to trade additional sampling steps for improved outcomes (Liu et al. 2023; Xu et al. 2024). Our goal is to generate high-quality samples within 2-4 steps to strike a favorable tradeoff between efficiency and efficacy. We observe that CMs enjoy the trade-off by alternating denoising and noise injection at inference time. However, CD approaches haven’t unleashed the potential power of the pre-trained teacher DMs, considering ODE solvers underperform SDE ones when the NFE is large. We first lay out a justification for the feasibility of combining CD with SDE solvers. Then, we identify several critical factors to make SCott workable for high-resolution text-to-image generation. We also introduce GAN loss to further enhance the sampling quality at rare steps. Our method is summarized in Figure 2.
@@
Figure 2: Overview of SCott. SCott distills a pre-trained teacher DM into a student one for accelerated sampling. Compared to the vanilla consistency distillation approach, we introduce a multi-step SDE solver to establish a stronger and more versatile teacher. We train the student model with CD loss using SDE solvers. Additionally, we include an adversarial learning loss to correct student output, boosting the sample quality with rare sampling steps. Note that we omit the EMA operation for the teacher for brevity.
Justification of Using SDE Solvers for CD
@@
Despite our motivation for combining SDE solvers and CD being sensible from the perspective of model distillation, a major concern arises as to whether it is reasonable to do so because CD is originally defined on ODE trajectories. In this section, we provide theoretical justification for this. Following (Song et al. 2023), we provide the convergence proof of using SDE solvers in CD.
Theorem 1.
Let $\Delta t:=\max\limits_{n\in[1,N]}{|t_{n+1}-t_{n}|}$ where $t\in[\tau,T]$. Assume ${\bm{f}}_{\theta}(\cdot,\cdot)$ is Lipschitz in ${\bm{x}}$ with constant $L_{1}$. Denote ${\bm{f}}(\cdot,\cdot)$ the consistency function of the SDE defined in Equation 3. Assume the SDE solver $\Phi_{SDE}$ has a local error bound of $O((\Delta t))^{p+1}$ with $p\geq 1$.Then, if $\mathcal{L}_{CD}^{N}({\bm{\theta}},\Phi_{SDE})=0$, we have :
$\sup\limits_{n,{\bm{x}}}||{\bm{f}}_{\theta}({\bm{x}}_{t_{n}},t_{n})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n})||_{2}=O((\Delta t)^{p})$
Proof.
We refer to Appendix 1 in our supplementary materials for the full proof.
Additionally, we also present an empirical investigation to demonstrate the feasibility of using SDE solvers for CD (Please refer to Appendix 3).
Stochastic Consistency Distillation (SCott)
@@
Controlling the Level of Noise in SDE Solvers
It is a natural idea to control the level of random noise injected in SDE solvers to stabilize training. Taking the DDIM solver in Equation 2 for example, the $\sigma_{t_{n}}$ controls the intensity of injected noise and typically we can introduce a coefficient $\eta$ for scaling it (Song, Meng, and Ermon 2020):
$\sigma_{t_{n}}(\eta):=\eta\sqrt{(1-\alpha_{t_{n-1}})/(1-\alpha_{t_{n}})}\sqrt{1-\alpha_{t_{n}}/\alpha_{t_{n-1}}}.$ (5)
Increasing $\eta$ from 0 to a positive value results in a set of SDE solvers with increasingly intensive noise, starting from an ODE one. We empirically experiment with the solvers for SCott and present the results in Table 5. As shown, we indeed need to select a reasonable noise level to enjoy the benefits of SDE solvers for error correction while avoiding introducing excessive variance. Following this insight, we include the more advanced ER-SDE solver (Cui et al. 2023) to SCott. The novel noise scaling function in the ER-SDE solver also provides a simple way to control the intensity of noise, which aligns with our requirements for SDE solvers.
Multi-step Sampling
The discretization errors of SDE solvers are larger than those of ODE solvers (Xu et al. 2023). A simple remediation to this is to decrease the step size, but doing so leads to slow convergence and degraded results for SCott, consistent with the results in (Luo et al. 2023a). To address this, we propose a multi-step sampling strategy for SCott. Specifically, for a sample ${\bm{z}}_{t_{n}}$ at the time $t_{n}$ on the SDE trajectory, to obtain the estimated state $\hat{{\bm{z}}}_{t_{m}}$ for a preceding timestep $t_{m}$, we split the time interval $t_{n}-t_{m}$ into $h$ intervals, namely, sampling with a step size of $\frac{t_{n}-t_{m}}{h}$, as illustrated in Figure 3. We denote the solution as $\hat{{\bm{z}}}_{t_{m}}=\Phi_{SDE}({\bm{\epsilon}}_{\theta},{\bm{z}}_{t_{n}},t_{n},t_{m},h)$.
Figure 3: Multi-step SDE solver sampling
@@
Introducing GAN into SCott
For high-resolution text-to-image generation, considering the high data dimensionality and complex data distribution, simply using L2 loss fails to capture data discrepancy precisely, thus providing imperfect consistency constraints. Suggested by CTM (Kim et al. 2023) which leverages GAN to improve trajectory estimation in the process of distillation, we integrate GAN into SCott to mitigate the restriction of L2 loss. However, it is non-trivial to directly employ CTM’s GAN. Firstly, CTM’s GAN performs on the pixel level while SCott performs on the latent space. For high-resolution image generation, employing a GAN that operates directly on pixel values would significantly increase the training cost. Secondly, SCott is designed to accelerate text-to-image tasks, the interaction between textual information and visual representation is not considered in the design of CTM’s GAN. To address these issues, we design a low-rank adaptation (LoRA) (Hu et al. 2021) discriminator with time and text projection, which renders adversarial training successfully in SCott.
LoRA Discriminator
@@
To accelerate training, the discriminator is initialized by the pre-trained U-Net in SD1.5. However, we find training the whole U-Net is expensive and unstable. Therefore, we freeze the encoder of the U-Net and only train the LoRA parameters of the decoder of the network. The frozen encoder of the U-Net extracts rich latent representations from the input and conditions, while the LoRA decoder acts as discriminator heads to distinguish between real images and generated images. By employing this approach, only $3.6\%$ of the parameters in the U-Net are updated, which not only contributes to computational efficiency but also increases the overall stability of the adversarial training.
Conditional Discriminator
Since SCott predicts the clean image $\hat{{\bm{z}}}_{0}$, we can directly perform adversarial learning between real image ${\bm{z}}_{0}$ and generated image $\hat{{\bm{z}}}_{0}$. However, we find such a naive design makes training unstable for text-to-image tasks. To surmount this challenge, the timestep $t_{n}$ and text ${\bm{c}}$ are propagated into the discriminator as conditions.
Following (Sauer et al. 2023a), we train our discriminator $D_{{\bm{\phi}}}$ with hinge loss:
$\begin{split}&\mathcal{L}_{adv}({\bm{\theta}},{\bm{\phi}})=\mathbb{E}_{{\bm{z}}_{0},{\bm{c}}}\Big{[}\max(0,1-D_{{\bm{\phi}}}({\bm{z}}_{0},{\bm{c}},0))\Big{]}\\ &\quad+\mathbb{E}_{{\bm{z}}_{t_{n}},{\bm{c}},t_{n}}\Big{[}\max(0,1+D_{{\bm{\phi}}}({\bm{f}}_{{\bm{\theta}}}({\bm{z}}_{t_{n}},{\bm{c}},\omega,t_{n}),{\bm{c}},t_{n}))\Big{]},\end{split}$ (6)
@@
where ${\bm{f}}_{{\bm{\theta}}}$ denotes the SCott model while $D_{{\bm{\phi}}}$ represents our proposed LoRA Discriminator. Overall, SCott is trained with the following objective:
$\begin{split}\mathcal{L}_{SCott}({\bm{\theta}},{\bm{\phi}})=\mathcal{L}_{CD}({\bm{\theta}},{\bm{\theta}}^{-};\Phi_{SDE})+\lambda_{adv}\mathcal{L}_{adv}({\bm{\theta}},{\bm{\phi}}).\end{split}$ (7)
@@
In practice, we set $\lambda_{adv}=0.4$ to control the strength of the discriminator for refining the outputs of ${\bm{f}}_{\bm{\theta}}$. To save GPU memory, the training is performed in latent space. We find the introduction of GAN loss renders SCott to produce more realistic outputs, particularly at rare steps. We present more details of the training process for SCott in Appendix 4.
Method References Step Time (s) FID $\downarrow$ CS $\uparrow$ CR $\uparrow$
DPM++ (Lu et al. 2022b) ArXiv (2022) 25 0.88 20.1 0.318 0.9118
DDIM (Song, Meng, and Ermon 2020) ICLR (2021) 50 -- 20.3 0.318 0.9130
PD (Salimans and Ho 2022) ICLR (2022) 4 0.21 26.4 0.300 --
CAD (Meng et al. 2023) CVPR (2023) 8 0.34 24.2 0.300 --
LCM (Luo et al. 2023a) ArXiv (2023) 2 0.10 30.4 0.293 0.7882
INSTAFLOW-0.9B (Liu et al. 2023) ICLR (2024) 1 0.09 23.4 0.304 --
INSTAFLOW-1.7B (Liu et al. 2023) ICLR (2024) 1 0.12 22.4 0.309 --
UFOGEN (Xu et al. 2024) CVPR (2024) 1 0.09 22.5 0.311 --
UFOGEN (Xu et al. 2024) CVPR (2024) 4 -- 22.1 0.307 --
SCott (Ours) 1 0.09 26.8 0.295 0.8254
SCott (Ours) 2 0.12 21.9 0.310 0.9114
SCott (Ours) 4 0.18 21.8 0.311 0.9145
@@
Table 2: Comparisons with the state-of-the-art methods on MSCOCO-2017 5K in terms of FID, CS, CR with backbone SD1.5.
Experiments
In this section, we elaborate on the experimental results of our proposed SCott model for the text-to-image generation task. We start with the comparison with other works, and then ablate components of SCott, highlighting the effectiveness of our proposed components.
Implementation Details
We use LAION-Aesthetics-6+ dataset (Schuhmann et al. 2022). We train SCott with 4 A100 GPUs and a batch size of 40 for 40K iterations. The learning rate is 8e-6 for SCott and 2e-5 for the discriminator. (See more experimental setups in Appendix 4 )
Method Step FID $\downarrow$
SWIFTBRUSH (Nguyen and Tran 2024) 1 16.67
UFOGEN (Xu et al. 2024) 1 12.78
INSTAFLOW-0.9B (Xu et al. 2024) 1 13.10
DMD (Yin et al. 2024) 1 11.49
SCott (Ours) 2 11.13
SCott (Ours) 4 10.68